if 2,-1,-3 are the eigen values of the matrix A then find the eigen values of the matrix A Inverse
1
Expert's answer
2020-12-01T06:23:31-0500

We claim that the eigenvalues of "A^{-1}" are "\frac{1}{2},-1,-\frac{1}{3}" . We will prove that in 2 steps :

  • Suppose that "v_\lambda\neq 0" is an eigenvector of "A" associated to an eigenvalue "\lambda" (where "\lambda \in \{2, -1, 3\}"). Therefore we have :

"\frac{1}{\lambda} v_\lambda = \frac{1}{\lambda} id \cdot v_\lambda = \frac{1}{\lambda} A^{-1}Av_\lambda = \lambda \cdot \frac{1}{\lambda}(A^{-1}v_\lambda) = A^{-1} v_\lambda"

So "v_\lambda" is an eigenvector of "A^{-1}" associated to an eigenvalue "\frac{1}{\lambda}" .

  • Now suppose that "v_\alpha\neq 0" is an eigenvector "A^{-1}" associated to an eigenvalue "\alpha" . We have :

"A(v_\alpha)=A (\frac{1}{\alpha}A^{-1}v_\alpha) = \frac{1}{\alpha} (A A^{-1}v_\alpha) = \frac{1}{\alpha} v_\alpha"

So we have that "\frac{1}{\alpha}" is an eigenvalue of "A" .

Thus we have proven that eigenvalues of "A^{-1}" are "\frac{1}{2}, -1, -\frac{1}{3}" and there is no other eigenvalues.